# https://github.com/OpenGVLab/VideoMamba VideoMamba: State Space Model for Efficient Video Understanding

### 11 Mar 2024 v1

@@ Abstract
†† Interns at Shanghai AI Laboratory. Corresponding author.Addressing the dual challenges of local redundancy and global dependencies in video understanding, this work innovatively adapts the Mamba to the video domain. The proposed VideoMamba overcomes the limitations of existing 3D convolution neural networks and video transformers. Its linear-complexity operator enables efficient long-term modeling, which is crucial for high-resolution long video understanding. Extensive evaluations reveal VideoMamba’s four core abilities: (1) Scalability in the visual domain without extensive dataset pretraining, thanks to a novel self-distillation technique; (2) Sensitivity for recognizing short-term actions even with fine-grained motion differences; (3) Superiority in long-term video understanding, showcasing significant advancements over traditional feature-based models; and (4) Compatibility with other modalities, demonstrating robustness in multi-modal contexts. Through these distinct advantages, VideoMamba sets a new benchmark for video understanding, offering a scalable and efficient solution for comprehensive video understanding. All the code and models are available.
@@ 2.2 Video Understanding
As for the architecture, it has evolved from using CNN which extracts features from video frames, to more advanced techniques. Initially, 3D CNNs [77, 10, 18, 78] expanded the traditional 2D CNN architecture to capture videos’ spatio-temporal information. Two-Stream [66], which combines spatial and temporal streams, and SlowFast [20], which uses parallel networks to capture semantics and rapid movements, further enhance action recognition capacity. The introduction of attention-based models [4, 2, 60, 7, 64], like TimeSformer [4] and ViViT [2], marked a significant advancement by effectively capturing long-range dependencies within video sequences, enhancing temporal relationship understanding. Recent developments [45, 43, 83, 53] have focused on accurate video transformer, with innovations like the VideoSwin’s window attention [53] and the UniFormer’s integration of convolution and self-attention mechanisms [45], aiming to balance computational efficiency with performance. Despite these models’ achievements in various tasks, they often come with high computational costs for long sequences. In contrast, our VideoMamba introduces a linear-complexity operator for efficient long-term modeling, outperforming existing methods with faster speed and lower GPU consumption.

### 12 Mar 2024 v2

@@ Abstract
†† Interns at Shanghai AI Laboratory. Corresponding authors.Addressing the dual challenges of local redundancy and global dependencies in video understanding, this work innovatively adapts the Mamba to the video domain. The proposed VideoMamba overcomes the limitations of existing 3D convolution neural networks and video transformers. Its linear-complexity operator enables efficient long-term modeling, which is crucial for high-resolution long video understanding. Extensive evaluations reveal VideoMamba’s four core abilities: (1) Scalability in the visual domain without extensive dataset pretraining, thanks to a novel self-distillation technique; (2) Sensitivity for recognizing short-term actions even with fine-grained motion differences; (3) Superiority in long-term video understanding, showcasing significant advancements over traditional feature-based models; and (4) Compatibility with other modalities, demonstrating robustness in multi-modal contexts. Through these distinct advantages, VideoMamba sets a new benchmark for video understanding, offering a scalable and efficient solution for comprehensive video understanding. All the code and models are available.
@@ 2.2 Video Understanding
As for the architecture, it has evolved from using CNN which extracts features from video frames, to more advanced techniques. Initially, 3D CNNs [76, 9, 17, 77] expanded the traditional 2D CNN architecture to capture videos’ spatio-temporal information. Two-Stream [65], which combines spatial and temporal streams, TSN [80], which proposes sparse sampling, and SlowFast [19], which uses parallel networks to capture semantics and rapid movements, further enhance action recognition capacity. The introduction of attention-based models [4, 2, 59, 63, 89], like TimeSformer [4] and ViViT [2], marked a significant advancement by effectively capturing long-range dependencies within video sequences, enhancing temporal relationship understanding. Recent developments [44, 42, 83, 52] have focused on accurate video transformer, with innovations like the VideoSwin’s window attention [52] and the UniFormer’s integration of convolution and self-attention mechanisms [44], aiming to balance computational efficiency with performance. Despite these models’ achievements in various tasks, they often come with high computational costs for long sequences. In contrast, our VideoMamba introduces a linear-complexity operator for efficient long-term modeling, outperforming existing methods with faster speed and lower GPU consumption.
@@ 5 Conclusion
Appendix
Arch. | Model | iso. | Extra | Input | #Param | FLOPs | K400 | |
Data | Size | (M) | (G) | Top-1 | Top-5 | |||
Trans. | STAM [63] | ✓ | IN-21K | 64224 | 121 | 104011 | 79.2 | - |
TimeSformer-L [4] | ✓ | IN-21K | 96224 | 121 | 238031 | 80.7 | 94.7 | |
ViViT-L [2] | ✓ | IN-21K | 16224 | 311 | 399234 | 81.3 | 94.7 | |
Mformer-HR [59] | ✓ | IN-21K | 16336 | 311 | 959310 | 81.1 | 95.2 | |
SSM | VideoMamba-Ti | ✓ | IN-1K | 8224 | 7 | 934 | 76.9 | 92.9 |
VideoMamba-Ti | ✓ | IN-1K | 16224 | 7 | 1734 | 78.1 | 93.5 | |
VideoMamba-Ti | ✓ | IN-1K | 32224 | 7 | 3434 | 78.8 | 93.9 | |
VideoMamba-Ti | ✓ | IN-1K | 64224 | 7 | 6934 | 79.6 | 94.2 | |
VideoMamba-Ti | ✓ | IN-1K | 64384 | 7 | 20234 | 80.3 | 94.8 | |
VideoMamba-S | ✓ | IN-1K | 8224 | 26 | 3434 | 79.3 | 94.2 | |
VideoMamba-S | ✓ | IN-1K | 16224 | 26 | 6834 | 80.8 | 94.8 | |
VideoMamba-S | ✓ | IN-1K | 32224 | 26 | 13534 | 81.5 | 95.2 | |
VideoMamba-S | ✓ | IN-1K | 64224 | 26 | 27134 | 81.8 | 95.3 | |
VideoMamba-S | ✓ | IN-1K | 64384 | 26 | 39534 | 82.7 | 95.6 | |
VideoMamba-M | ✓ | IN-1K | 8224 | 74 | 10134 | 80.6 | 94.6 | |
VideoMamba-M | ✓ | IN-1K | 16224 | 74 | 20234 | 81.9 | 95.4 | |
VideoMamba-M | ✓ | IN-1K | 32224 | 74 | 40334 | 82.4 | 95.7 | |
VideoMamba-M | ✓ | IN-1K | 64224 | 74 | 80634 | 82.8 | 96.0 | |
VideoMamba-M | ✓ | IN-1K | 64384 | 74 | 236834 | 83.3 | 96.1 | |
Trans. | ST-MAE-B [18] | ✓ | 16224 | 87 | 18037 | 81.3 | 94.9 | |
VideoMAE-S [74] | ✓ | 16224 | 22 | 5735 | 79.0 | 93.8 | ||
VideoMAE-B [74] | ✓ | 16224 | 87 | 18035 | 81.5 | 95.1 | ||
UMT-B [43] | ✓ | CLIP-400M | 8224 | 87 | 18035 | 85.7 | 97.0 | |
SSM | VideoMamba-M | ✓ | CLIP-400M | 8224 | 74 | 10134 | 82.0 | 95.4 |
VideoMamba-M | ✓ | CLIP-400M | 16224 | 74 | 20234 | 83.4 | 95.9 | |
VideoMamba-M | ✓ | CLIP-400M | 32224 | 74 | 40334 | 83.9 | 96.2 | |
VideoMamba-M | ✓ | CLIP-400M | 64224 | 74 | 80634 | 84.3 | 96.6 | |
VideoMamba-M | ✓ | CLIP-400M | 64384 | 74 | 236834 | 85.0 | 96.9 | |
Appendix 0.A More Results
In Table IX, we present additional results on the Kinetics-400 dataset. These results clearly demonstrate that our SSM-based model outperforms all previous attention-based methods. We observe consistent performance improvements with increasing resolution and frame count.
config | Single-Modality | Multi-Modality | |
SthSthV2 | K400 | 5M & 17M & 25M | |
optimizer | AdamW | AdamW | |
optimizer momentum | |||
weight decay | 0.05 | 0.05 | |
learning rate schedule | cosine decay | cosine decay | |
learning rate | 1.2e-3 | 4e-4 | |
minimal learning rate | 1e-5 | 4e-6 | |
batch size | 2048 | 2048 , 2048 | |
warmup epochs | 40 | 1 | |
total epochs | 800 | 10 | |
mask ratio | 80% | 50% , 80% , 50% | |
input frame | 8 | 8 | |
drop path | 0.4 | 0.25 | |
flip augmentation | no | yes | yes |
augmentation | MultiScaleCrop[0.66, 0.75, 0.875, 1] | MultiScaleCrop[0.5, 1] |
config | 224224 | 448448 | 512512 |
optimizer | AdamW | ||
optimizer momentum | |||
weight decay | 0.1(Ti), 0.05(S,M) | 1e-8 | 1e-8 |
learning rate schedule | cosine decay | ||
base learning rate | 5e-4 | 5e-6 | 5e-6 |
minimal learning rate | 1e-5 | 5e-6 | 5e-6 |
base batch size | 512 | ||
repeated augmentation | no(Ti), yes(S,M) | ||
warmup epochs | 5(Ti,S), 30(M) | 5 | 2 |
total epochs | 300 | 30 | 10 |
drop path | 0(Ti), 0.15(S), 0.5(M) | ||
label smoothing | 0.1 | ||
cutmix | 1.0 | ||
augmentation | RandAug(7, 0.25)(Ti), RandAug(9, 0.5)(S,M) |
config | 224224 | 384384 |
optimizer | AdamW | |
optimizer momentum | ||
weight decay | 0.1(Ti), 0.05(S,M,M) | 1e-8 |
learning rate schedule | cosine decay | |
base learning rate | 4e-4(Ti,S), 2e-4(M), 1e-4(M) | 5e-6 |
minimal learning rate | 1e-6 | |
base batch size | 256 | |
repeated augmentation | 2 | |
warmup epochs | 5 | 2 |
total epochs | 70(Ti), 50(S,M), 45(M) | 10 |
drop path | 0.1(Ti), 0.35(S), 0.8(M), 0.4(M) | |
layer-wise lr decay | 0.75(S,M,M), 0.8(M) | |
flip augmentation | yes | |
label smoothing | 0.1 | |
cutmix | 1.0 | |
augmentation | RandAug(7, 0.25)(Ti), RandAug(9, 0.5)(S,M,M) |
config | 224224 | 288288 |
optimizer | AdamW | |
optimizer momentum | ||
weight decay | 0.1(Ti), 0.05(S,M,M) | 1e-8 |
learning rate schedule | cosine decay | |
base learning rate | 4e-4(Ti,S,M) 1e-4(M) | 5e-6 |
minimal learning rate | 1e-6 | |
base batch size | 256 | |
repeated augmentation | 2 | |
warmup epochs | 5 | 2 |
total epochs | 35(Ti), 30(S,M,M) | 10 |
drop path | 0.1(Ti), 0.35(S), 0.8(M), 0.4(M) | |
layer-wise lr decay | 0.75(S,M,M), 0.8(M) | |
flip augmentation | no | |
label smoothing | 0.1 | |
cutmix | 1.0 | |
augmentation | RandAug(7, 0.25)(Ti), RandAug(9, 0.5)(S,M,M) |
config | BreakFast & LVU | COIN |
optimizer | AdamW | |
optimizer momentum | ||
weight decay | 0.1(Ti), 0.05(S,M,M) | |
learning rate schedule | cosine decay | |
base learning rate | 2e-4 | |
minimal learning rate | 1e-6 | |
base batch size | 256 | |
repeated augmentation | 2 | |
warmup epochs | 5 | |
total epochs | 70(Ti), 50(S,M), 45(M) | 40(Ti), 35(S), 30(M,M) |
drop path | 0.1(Ti), 0.35(S), 0.8(M), 0.4(M) | |
layer-wise lr decay | 0.75(S,M,M), 0.8(M) | |
flip augmentation | yes | |
label smoothing | 0.1 | |
cutmix | 1.0 | |
augmentation | RandAug(7, 0.25)(Ti), RandAug(9, 0.5)(S,M,M) |
Appendix 0.B More Implementation Details
0.B.1 Training Details
We sparsely sample frames from the raw videos as in TSN [80] for all the datasets. Table X details the masked pretraining hyperparameters. For the unmasked multi-modality pretraining, we load the pretrained model and train it for an additional epoch with a learning rate of 8e-5. Moreover, Tables XI, XII, XIII, and XIV show the training details for the different datasets used for fine-tuning.
0.B.2 Dataset Descriptions
Dataset | #image/video | #text | Type |
COCO | 113K | 567K | image |
Visual Genome | 100K | 768K | image |
SBU Captions | 860K | 860K | image |
CC3M | 2.88M | 2.88M | image |
CC12M | 11.00M | 11.00M | image |
WebVid-2M | 2.49M | 2.49M | video |
WebVid-10M | 10.73M | 10.73M | video |
5M corpus = CC3MWebVid-2M | 5.37M | 5.37M | video+image |
17M corpus = 5MCOCOVGSBUCC12M | 17.44M | 18.57M | video+image |
25M corpus = 17MWebVid-10MWebVid-2M | 25.68M | 26.81M | video+image |
Dataset | #video | #text | Avg Video | ||||
Train | Val | Test | Train | Val | Test | Length (s) | |
Image Classification | |||||||
ImageNet-1K | 1,281,167 | 50,000 | 100,000 | - | - | - | - |
Short-term Action Recognition | |||||||
Kinetics-400 | 240,436 | 19,787 | - | - | - | - | 10 |
Something-Something V2 | 168,913 | 24,777 | - | - | - | - | 4 |
Long-term Action Recognition | |||||||
Breakfast | 1,577 | - | 410 | - | - | - | 137 |
COIN | 9,026 | - | 2,796 | - | - | - | 142 |
LVU | 7,619 | 1,666 | 1,551 | - | - | - | 134 |
Relation | 138 | 49 | 41 | - | - | - | 127 |
Speak | 871 | 196 | 188 | - | - | - | 133 |
Scene | 514 | 107 | 81 | - | - | - | 132 |
Director | 680 | 163 | 107 | - | - | - | 137 |
Genre | 2807 | 569 | 584 | - | - | - | 130 |
Writer | 748 | 174 | 168 | - | - | - | 142 |
Year | 725 | 163 | 141 | - | - | - | 133 |
Like | 658 | 142 | 139 | - | - | - | 159 |
View | 478 | 103 | 102 | - | - | - | 112 |
Video-Text Retrieval | |||||||
MSRVTT | 7,010 | - | 1,000 | 140,200 | - | 1,000 | 15 |
DiDeMo | 8,496 | 1,094 | 1,036 | 8,496 | 1,094 | 1,036 | 29 |
ActivityNet | 10,009 | 4,917 | - | 10,009 | 4,917 | - | 180 |
LSMDC | 101,055 | - | 1,000 | 101,055 | - | 1,000 | 5 |
MSVD | 1,200 | 100 | 670 | 1,200 | 100 | 670 | 15 |